OpenStudy (anonymous):

simplify completely 12x+36/x^2-4x-21

OpenStudy (anonymous):

12(x+3)/(x-7)(x+3) =12/(x-7)

OpenStudy (anonymous):

theres these symbols next to them with a -3 or a -7 or both \[\neq7\]

OpenStudy (anonymous):

thats what it looks like

OpenStudy (anonymous):

which ones right?

OpenStudy (anonymous):

@sunnyshores

OpenStudy (anonymous):

Pretty sure it's -7

OpenStudy (anonymous):

yea you had it right but next to it those numbers are there they dont have an option without it thhey look like this \[\frac{ 12 }{ x-7 } x \neq7\]

OpenStudy (anonymous):

@sunnyshores

OpenStudy (anonymous):

ohhhh so are u asking if is -7 or \[\neq7\] ?

OpenStudy (anonymous):

no im asking if it should have just the 7 or a minus 3 too

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thats a good question! let me look at the equation again ;)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

it should have the -7 and the -3 too :D u were right! :D

OpenStudy (anonymous):

thnk you :)

OpenStudy (anonymous):

np!